OpenStudy (anonymous):

write an equation of a circle with a center at (1,2) and a diameter of 15

OpenStudy (whpalmer4):

General formula for a circle with radius \(r\) and center at \((h,k)\) is \[(x-h)^2+(y-k)^2=r^2\]Fill in the blanks.

random231 (random231):

@whpalmer4 is right!

OpenStudy (anonymous):

well the equation of a circles is:\[(x-h)^2+(y-k)^2=r^2\]

OpenStudy (anonymous):

oh wow @whpalmer4 already said that

OpenStudy (anonymous):

lol

OpenStudy (anonymous):

in your case r = 15/2